# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ V IQ-Aware Hybrid Precoding
V-A IQ-Aware Hybrid Precoding with FC-PS Network
V-A1 Problem statement
We first consider the IQ-aware hybrid precoding with a FC-PS network, where each RF chain is connected to all BS antennas via PSs. Denote the number of RF chains as $N_{\rm RF}$, satisfying $N_{s}\leq N_{\rm RF}\ll N_{t}$. In the baseband processor, an IQ-aware digital precoder $\bar{{\mathbf{D}}}$ of dimension $2N_{\rm RF}\times 2N_{s}$ is employed, with block components ${\mathbf{D}}_{k\ell}\in\mathbb{R}^{N_{\rm RF}\times N_{s}},k\in\{1,2\},\ell\in\{1,2\}$, to precode the symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$. The complex signal vector coming out of RF chains is thus represented as ${\mathbf{s}}_{\rm rf}={\mathbf{D}}_{11}{\mathbf{s}}_{I}+{\mathbf{D}}_{12}{\mathbf{s}}_{Q}+j({\mathbf{D}}_{21}{\mathbf{s}}_{I}+{\mathbf{D}}_{22}{\mathbf{s}}_{Q})$. At the RF front-end, the analog precoder implemented by PSs is expressed by a complex matrix ${\mathbf{A}}={\mathbf{A}}_{I}+j{\mathbf{A}}_{Q}\in\mathbb{C}^{N_{t}\times N_{\rm RF}}$. Given that PSs can only adjust the phases of signals, each element of ${\mathbf{A}}$ must satisfy the unit modulus constraint: $|{\mathbf{A}}_{m,n}|^{2}={\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\forall m,n$. Thereafter, the transmitted signal ${\mathbf{x}}$ is given as ${\mathbf{x}}={\mathbf{A}}{\mathbf{s}}_{\rm rf}$, whose real form is
The covariance matrix of $\bar{{\mathbf{x}}}$ is thus $\bar{{\mathbf{Q}}}=\frac{1}{2}\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{A}}}^{T}$. Comparing precoding models (33) and (46), one can observe that the anlog precoder $\bar{{\mathbf{A}}}$ inherently possesses coupled diagonal and off-diagonal block components, which differentiates IQ-aware hybrid precoding from IQ-aware fully digital precoding.
To maximize the achievable rate, the analog precoder $\bar{{\mathbf{A}}}$ and digital precoder $\bar{{\mathbf{D}}}$ are jointly designed to approximate the optimal IQ-aware digital precoder $\bar{{\mathbf{F}}}$. Analogous to classical hybrid precoding design, we adopt the least-square principle to formulate the problem:
$(\mathcal{P}1)\quad\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (47a)
where (47b) is the total power constraint. Owing to the non-convex modulus constraint in (47a), the coupled components of $\bar{{\mathbf{A}}}$, as well as the product of $\bar{{\mathbf{A}}}$ and $\bar{{\mathbf{D}}}$, directly solving problem (47) is challenging.
@@ V-A2 Problem transformation
To overcome these challenges, our first step is to seek an equivalent expression of (47) that decouples the analog precoder $\bar{{\mathbf{A}}}$ from its multiplication with $\bar{{\mathbf{D}}}$, which can significantly simplify the precoding design.
Note that the columns of the optimal precoder $\bar{{\mathbf{F}}}$ are mutually orthogonal to mitigate the interference between data streams. The designed hybrid precoder $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$ is thus expected to maintain column orthogonality as well. In massive MIMO systems, $N_{t}\gg N_{RF}$, the relationship $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\approx N_{t}{\mathbf{I}}_{2N_{\rm RF}}$ holds with high probability [35]. This is because the diagonal elements of $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}$ are exactly $N_{t}$ while the off-diagonal elements can be approximated as the summation of $2N_{t}$ independent terms, which is much less than $N_{t}$ with high probability for $N_{t}\gg 1$. These properties imply that the columns of the optimal $\bar{{\mathbf{D}}}$ are approximately orthogonal as well. Motivated by this observation, we introduce a column-orthogonal constraint on $\bar{{\mathbf{D}}}$ and thereby replace it by $\gamma\bar{{\mathbf{D}}}_{u}$, where $\bar{{\mathbf{D}}}^{T}_{u}\bar{{\mathbf{D}}}_{u}={\mathbf{I}}_{2N_{s}}$ and $\gamma$ is a normalization scalar to satisfy the total power constraint (47b). To determine the value of $\gamma$, we can substitute $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\approx N_{t}{\mathbf{I}}_{2N_{\rm RF}}$ and $\gamma\bar{{\mathbf{D}}}_{u}$ into (47b), thereby obtaining that $\gamma\approx\sqrt{\frac{2P}{N_{t}}}$. Consequently, the hybrid precoding design problem is recast as
@@
$(\mathcal{P}2)\quad\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (48a)
@@
Given matrices $\bar{{\mathbf{F}}}\in\mathbb{R}^{2N_{t}\times 2N_{s}}$, $\bar{{\mathbf{A}}}\in\mathbb{R}^{2N_{t}\times 2N_{\rm RF}}$, and $\bar{{\mathbf{D}}}_{u}\in\mathbb{R}^{2N_{\rm RF}\times 2N_{s}}$ satisfying $\bar{{\mathbf{D}}}_{u}^{T}\bar{{\mathbf{D}}}_{u}={\mathbf{I}}_{2N_{s}}$ and $N_{t}\gg N_{RF}\geq N_{s}$, the following optimization problems are identical:
@@
Comparing (47) and (50), it becomes evident that the analog precoder $\bar{{\mathbf{A}}}$ no longer appears in the product form with the digital precoder $\widetilde{{\mathbf{D}}}$. This simplification significantly facilitates our precoding algorithm design. By applying Lemma 1 and (50), problem $(\mathcal{P}2)$ is equivalently transformed into
@@
$(\mathcal{P}3)\quad\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (51a)
@@
To efficiently solve this problem, we exploit the alternating minimization method, where the three variables $\bar{{\mathbf{A}}}$, $\widetilde{{{\mathbf{D}}}}$, and $\bar{{\mathbf{F}}}_{c}$ are updated alternatively while keeping the others fixed until convergence. If the updated solutions in each iteration are all optimal, the alternating minimization method is guaranteed to converge to a local optimum, as elaborated in Section V-C.
@@ V-A3 Algorithm design
Algorithm 1 : IQ-aware hybrid precoding with FC-PS network
@@
The step-by-step procedures of the proposed IQ-aware hybrid precoding algorithm for FC PSs are summarized in Algorithm 1, with detailed explanations provided below.
i) Fix ($\widetilde{{{\mathbf{D}}}}$, $\bar{{{\mathbf{F}}}}_{c}$) and optimize $\bar{{\mathbf{A}}}$: We start by deriving the optimal solution for $\bar{{\mathbf{A}}}$ after fixing matrices $\widetilde{{{\mathbf{D}}}}$ and $\bar{{{\mathbf{F}}}}_{c}$. We denote the block matrix components of $\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}$ as ${\mathbf{Z}}_{k\ell}\in\mathbb{R}^{N_{t}\times N_{\rm RF}}$, $k\in\{1,2\},\>\ell\in\{1,2\}$, such that
@@
Then, the objective function can be rewritten as
@@
where ${\mathbf{Z}}_{I}\overset{\Delta}{=}{\mathbf{Z}}_{11}+{\mathbf{Z}}_{22}$, ${\mathbf{Z}}_{Q}\overset{\Delta}{=}{\mathbf{Z}}_{21}-{\mathbf{Z}}_{12}$, and $C_{1}=\gamma^{2}\|{\mathbf{A}}_{I}\|_{F}^{2}+\gamma^{2}\|{\mathbf{A}}_{Q}\|_{F}^{2}-2\mathsf{Tr}({\mathbf{Z}}_{11}^{T}{\mathbf{Z}}_{22})+2\mathsf{Tr}({\mathbf{Z}}_{21}^{T}{\mathbf{Z}}_{12})=\gamma^{2}N_{t}N_{\rm RF}-2\mathsf{Tr}({\mathbf{Z}}_{11}^{T}{\mathbf{Z}}_{22})+2\mathsf{Tr}({\mathbf{Z}}_{21}^{T}{\mathbf{Z}}_{12})$. Equations (57) and (51a) imply that the joint optimization of analog precoder is equivalent to independently optimizing each element of ${\mathbf{A}}={\mathbf{A}}_{I}+j{\mathbf{A}}_{Q}$. Consider the PS connecting the $m$-th antenna and the $n$-th RF chain. The optimal precoder coefficient is obtained by solving
@@
$(\mathcal{P}4)\quad\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1.$ (59a)
Problem $(\mathcal{P}4)$ can be interpreted as finding a point $({\mathbf{A}}_{I,m,n},{\mathbf{A}}_{Q,m,n})$ on the unit circle that is closest to $(\frac{1}{\gamma}{\mathbf{Z}}_{I,m,n},\frac{1}{\gamma}{\mathbf{Z}}_{Q,m,n})$. The close-form solution is given as
$\left\{\begin{array}[]{c}{\mathbf{A}}^{\star}_{I,m,n}=\frac{{\mathbf{Z}}_{I,m,n}}{\sqrt{{\mathbf{Z}}_{I,m,n}^{2}+{\mathbf{Z}}_{Q,m,n}^{2}}}\\ {\mathbf{A}}_{Q,m,n}^{\star}=\frac{{\mathbf{Z}}_{Q,m,n}}{\sqrt{{\mathbf{Z}}_{I,m,n}^{2}+{\mathbf{Z}}_{Q,m,n}^{2}}}\end{array}\right.,$ (62)
which completes the update of $\bar{{\mathbf{A}}}$ in step 4 of Algorithm 1.
ii) Fix ($\bar{{{\mathbf{A}}}}$, $\bar{{{\mathbf{F}}}}_{c}$) and optimize $\widetilde{{\mathbf{D}}}$: With fixed matrices $\bar{{{\mathbf{A}}}}$ and $\bar{{{\mathbf{F}}}}_{c}$, the objective function for optimizing $\bar{{\mathbf{D}}}_{u}$ can be rewritten as
@@
where (a) holds because $\widetilde{{\mathbf{D}}}^{T}\widetilde{{\mathbf{D}}}={\mathbf{I}}_{2N_{\rm RF}}$ and $\mathsf{Tr}\{\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\}=2N_{t}N_{\rm RF}$. Thereafter, the optimal digital precoder $\widetilde{{\mathbf{D}}}$ is updated by solving
$\max_{\widetilde{{{\mathbf{D}}}}}\>\>$ $\mathsf{Tr}\{\widetilde{{\mathbf{D}}}\widetilde{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}\}$ (64)
$(\mathcal{P}5)\quad\mathrm{s.t.}\>\>$ $\widetilde{{{\mathbf{D}}}}^{T}\widetilde{{\mathbf{D}}}={\mathbf{I}}_{2N_{\rm RF}}.$ (64a)
Problem $(\mathcal{P}5)$ is a standard Orthogonal Procrustes Problem (OPP). To solve it, we compute the SVD of $\widetilde{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}$ to obtain ${\mathbf{V}}_{1}\bm{\Sigma}{\mathbf{V}}_{2}^{T}$. Then, (64) has the following upper bound:
@@
where (b) arises because the matrix ${\mathbf{V}}_{2}^{T}\widetilde{{\mathbf{D}}}{\mathbf{V}}_{1}$ is unitary. The equality holds if and only if ${\mathbf{V}}_{2}^{T}\widetilde{{\mathbf{D}}}{\mathbf{V}}_{1}={\mathbf{I}}_{2N_{\rm RF}}$. Thereafter, $\widetilde{{\mathbf{D}}}$ is updated by
$\widetilde{{\mathbf{D}}}^{\star}={\mathbf{V}}_{2}{\mathbf{V}}_{1}^{T}.$ (66)
Accordingly, the optimal digital precoder $\bar{{\mathbf{D}}}_{u}^{\star}$ is determined by the first $2N_{s}$ columns of $\widetilde{{\mathbf{D}}}^{\star}$:
$\bar{{\mathbf{D}}}_{u}^{\star}=\widetilde{{\mathbf{D}}}^{\star}_{:,1:2N_{s}},$ (67)
which completes the update of digital precoder $\bar{{\mathbf{D}}}_{u}$ in steps 5 and 6.
@@
where $C_{2}=\|\bar{{\mathbf{F}}}-\gamma\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{u}\|_{F}^{2}$ is irrelevant to $\bar{{\mathbf{F}}}_{c}$. It is evident that the optimal $\bar{{\mathbf{F}}}_{c}$ needs to cancel the least square error introduced by $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{c}$:
@@
iv) Power normalization: Recall that the previously calculated $\gamma$ is an approximated value. After the alternating minimization algorithm converges, we normalize the digital precoder $\bar{{\mathbf{D}}}_{u}$ by $\sqrt{\frac{2P}{\mathsf{Tr}\{\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{u}\bar{{\mathbf{D}}}^{T}_{u}\bar{{\mathbf{A}}}^{T}\}}}\bar{{\mathbf{D}}}_{u}$ to guarantee that the optimized precoders can perfectly satisfy the total power constraint. This completes the design of IQ-aware hybrid precoding with FC-PS structure.
V-B IQ-Aware Hybrid Precoding with SC-PS Network

### v2

@@ V-A2 Problem transformation
Owing to the non-convex modulus constraint in (43a), the block-coupled constraint of $\bar{{\mathbf{A}}}$, as well as the product of $\bar{{\mathbf{A}}}$ and $\bar{{\mathbf{D}}}$, directly solving problem (43) is challenging. To overcome these challenges, the first step is to seek an equivalent expression of (43) that decouples the analog precoder $\bar{{\mathbf{A}}}$ from its multiplication with $\bar{{\mathbf{D}}}$, which can significantly simplify the precoding design.
Note that the columns of the optimal precoder $\bar{{\mathbf{F}}}$ are mutually orthogonal to mitigate the interference between data streams. The designed hybrid precoder $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$ is thus expected to maintain column orthogonality as well. In massive MIMO systems, $N_{t}\gg N_{RF}$, the relationship $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\approx N_{t}{\mathbf{I}}_{2N_{\rm RF}}$ holds with high probability [37, 38]. This is because the diagonal elements of $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}$ are exactly $N_{t}$ while the off-diagonal elements can be approximated as the summation of $2N_{t}$ independent terms, which is much less than $N_{t}$ with high probability for $N_{t}\gg 1$. These properties imply that the columns of the optimal $\bar{{\mathbf{D}}}$ are approximately orthogonal as well. Motivated by this observation, we introduce a column-orthogonal constraint on $\bar{{\mathbf{D}}}$ and thereby replace it by $\gamma\bar{{\mathbf{D}}}_{u}$, where $\bar{{\mathbf{D}}}^{T}_{u}\bar{{\mathbf{D}}}_{u}={\mathbf{I}}_{2N_{s}}$. The scalar $\gamma$ is introduced to meet the total power constraint (43b), whose value is approximately $\gamma\approx\sqrt{\frac{2P}{N_{t}}}$ by considering $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\approx N_{t}{\mathbf{I}}_{2N_{\rm RF}}$. Consequently, the hybrid precoding design is recast as
@@
$\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (44a)
@@
Given matrices $\bar{{\mathbf{F}}}\in\mathbb{R}^{2N_{t}\times 2N_{s}}$, $\bar{{\mathbf{A}}}\in\mathbb{R}^{2N_{t}\times 2N_{\rm RF}}$, and $\bar{{\mathbf{D}}}_{u}\in\mathbb{R}^{2N_{\rm RF}\times 2N_{s}}$ satisfying $\bar{{\mathbf{D}}}_{u}^{T}\bar{{\mathbf{D}}}_{u}={\mathbf{I}}_{2N_{s}}$ and $N_{RF}\geq N_{s}$, the following optimization problems are identical:
@@
Equation (46) gets the analog precoder $\bar{{\mathbf{A}}}$ rid of the product form with the digital precoder $\widetilde{{\mathbf{D}}}$. This simplification significantly facilitates our precoding algorithm design. By applying Lemma 3 and (46), problem (44) is transformed into
@@
$\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (47a)
@@
To efficiently solve this problem, we exploit the alternating minimization method. Specifically, the three variables $\bar{{\mathbf{A}}}$, $\widetilde{{{\mathbf{D}}}}$, and $\bar{{\mathbf{F}}}_{c}$ are updated alternatively while keeping the others fixed until convergence. If the updated solutions in each iteration are all optimal, the alternating minimization method is guaranteed to converge to a local optimum.
@@ V-A3 Algorithm design
Algorithm 1 : IQ-aware hybrid precoding with a FC-PS network
@@
The step-by-step procedures of the proposed IQ-aware hybrid precoding algorithm for a FC-PS network are summarized in Algorithm 1, with detailed explanations provided below.
i) Fix ($\widetilde{{{\mathbf{D}}}}$, $\bar{{{\mathbf{F}}}}_{c}$) and optimize $\bar{{\mathbf{A}}}$: We start by deriving the optimal solution for $\bar{{\mathbf{A}}}$ after fixing matrices $\widetilde{{{\mathbf{D}}}}$ and $\bar{{{\mathbf{F}}}}_{c}$. Denote the block matrix components of $\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}$ as ${\mathbf{Z}}_{k\ell}\in\mathbb{R}^{N_{t}\times N_{\rm RF}}$, $k\in\{1,2\},\>\ell\in\{1,2\}$, such that
@@
By substituting (50) into (47), the objective function can be rewritten as
@@
wherein ${\mathbf{Z}}_{I}\overset{\Delta}{=}{\mathbf{Z}}_{11}+{\mathbf{Z}}_{22}$, ${\mathbf{Z}}_{Q}\overset{\Delta}{=}{\mathbf{Z}}_{21}-{\mathbf{Z}}_{12}$, and $C_{1}=\gamma^{2}\|{\mathbf{A}}_{I}\|_{F}^{2}+\gamma^{2}\|{\mathbf{A}}_{Q}\|_{F}^{2}-2\mathsf{Tr}({\mathbf{Z}}_{11}^{T}{\mathbf{Z}}_{22})+2\mathsf{Tr}({\mathbf{Z}}_{21}^{T}{\mathbf{Z}}_{12})=\gamma^{2}N_{t}N_{\rm RF}-2\mathsf{Tr}({\mathbf{Z}}_{11}^{T}{\mathbf{Z}}_{22})+2\mathsf{Tr}({\mathbf{Z}}_{21}^{T}{\mathbf{Z}}_{12})$. Equations (53) and (47a) indicate that the joint optimization of analog precoder, ${\mathbf{A}}$, is equivalent to independently optimizing its elements. Consider the PS connecting the $m$-th antenna and the $n$-th RF chain. The optimal precoder coefficient is obtained by solving
@@
$\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1.$ (55a)
Problem (55) can be interpreted as finding a point $({\mathbf{A}}_{I,m,n},{\mathbf{A}}_{Q,m,n})$ on the unit circle that is closest to $(\frac{1}{\gamma}{\mathbf{Z}}_{I,m,n},\frac{1}{\gamma}{\mathbf{Z}}_{Q,m,n})$, whose optimal solution is
$\left\{\begin{array}[]{c}{\mathbf{A}}^{\star}_{I,m,n}=\frac{{\mathbf{Z}}_{I,m,n}}{\sqrt{{\mathbf{Z}}_{I,m,n}^{2}+{\mathbf{Z}}_{Q,m,n}^{2}}}\\ {\mathbf{A}}_{Q,m,n}^{\star}=\frac{{\mathbf{Z}}_{Q,m,n}}{\sqrt{{\mathbf{Z}}_{I,m,n}^{2}+{\mathbf{Z}}_{Q,m,n}^{2}}}\end{array}\right..$ (58)
This completes the update of ${\mathbf{A}}$ in Algorithm 1.
ii) Fix ($\bar{{{\mathbf{A}}}}$, $\bar{{{\mathbf{F}}}}_{c}$) and optimize $\widetilde{{\mathbf{D}}}$: With fixed matrices $\bar{{{\mathbf{A}}}}$ and $\bar{{{\mathbf{F}}}}_{c}$, the objective function in (47) can be rewritten as
@@
where (a) holds because $\widetilde{{\mathbf{D}}}^{T}\widetilde{{\mathbf{D}}}={\mathbf{I}}_{2N_{\rm RF}}$ and $\mathsf{Tr}\{\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}\}=2N_{t}N_{\rm RF}$. Accordingly, the sub-problem for optimizing $\widetilde{{\mathbf{D}}}$ is equivalent to
$\max_{\widetilde{{{\mathbf{D}}}}^{T}\widetilde{{\mathbf{D}}}={\mathbf{I}}_{2N_{\rm RF}}}\>\>$ $\mathsf{Tr}\{\widetilde{{\mathbf{D}}}\widetilde{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}\}$ (60)
Problem (60) is an orthogonal Procrustes Problem. To solve it, we can compute the SVD: $\widetilde{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}={\mathbf{V}}_{1}\bm{\Sigma}{\mathbf{V}}_{2}^{T}$. Then, the objective function in (60) is upper bounded by
@@
where (b) arises because ${\mathbf{V}}_{2}^{T}\widetilde{{\mathbf{D}}}{\mathbf{V}}_{1}$ is an orthogonal matrix. The equality holds if and only if ${\mathbf{V}}_{2}^{T}\widetilde{{\mathbf{D}}}{\mathbf{V}}_{1}={\mathbf{I}}_{2N_{\rm RF}}$. Henceforth, the optimal $\widetilde{{\mathbf{D}}}$ is obtained as
$\widetilde{{\mathbf{D}}}^{\star}={\mathbf{V}}_{2}{\mathbf{V}}_{1}^{T},$ (62)
which allows us to update $\bar{{\mathbf{D}}}_{u}^{\star}$ by $\widetilde{{\mathbf{D}}}^{\star}_{:,1:2N_{s}}$ in step 6 of Algorithm 1.
@@
where $C_{2}=\|\bar{{\mathbf{F}}}-\gamma\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{u}\|_{F}^{2}$ is irrelevant to $\bar{{\mathbf{F}}}_{c}$. It is evident that the optimal $\bar{{\mathbf{F}}}_{c}$ needs to cancel the least square error introduced by $\gamma\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{c}$:
@@
iv) Power normalization: Recall that the previously calculated $\gamma$ is an approximated value. After the alternating minimization algorithm converges, we normalize the digital precoder $\bar{{\mathbf{D}}}_{u}$ by $\sqrt{\frac{2P}{\mathsf{Tr}\{\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}_{u}\bar{{\mathbf{D}}}^{T}_{u}\bar{{\mathbf{A}}}^{T}\}}}\bar{{\mathbf{D}}}_{u}$ to guarantee the total power constraint. This completes the design of IQ-aware hybrid precoding with a FC-PS network.
V-B IQ-Aware Hybrid Precoding with a SC-PS Network